OpenStudy (anonymous):

Fan and Medal The change in water level of one portion of the mississippi River is about -0.3 ft per day. If the water level starts at 17 feet and falls for x days, the level represented by the function f(x) = 17 - 0.3x a. Find the intercepts and graph the function b. What does each intercept represent

OpenStudy (michele_laino):

please if I set x=0 into the formula for f(x), what do you get?

OpenStudy (anonymous):

What?

OpenStudy (michele_laino):

please what is f(0), namely what is 17-0.3*0=?

OpenStudy (anonymous):

It Equals 17

OpenStudy (michele_laino):

ok! y=17 is your y-intercept

OpenStudy (michele_laino):

Now, please if I set f(x)=0, I have thisequation below: -0.3*x+17=0 please compute x

OpenStudy (anonymous):

I dont get what your trying to say

OpenStudy (michele_laino):

please solve this equation: -0.3*x+17=0

OpenStudy (anonymous):

x=56.666667

OpenStudy (michele_laino):

that's right! So x=56.67 (approximated value) or x=170/3 is your x-intercept

OpenStudy (anonymous):

And how do i Graph that

OpenStudy (michele_laino):

it is vey simple, since your line has to pass at point (0,17) and it has to pass at point (56.67,0)

OpenStudy (michele_laino):



OpenStudy (michele_laino):

oops..it is very simple...

OpenStudy (anonymous):

Its okay. So if im right the bottom half gets shaded in?

OpenStudy (michele_laino):

no, I don't think, because your line represents the function f(x), of course you have to consider positive values of x only, and positive values of f(x) only

OpenStudy (michele_laino):